OpenStudy (anonymous):

Find the slope of a line that passes through (–2, –3), (1, 1).

OpenStudy (anonymous):

Do you know the slope formula?

OpenStudy (anonymous):

ehh no:/

OpenStudy (anonymous):

slope = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

If x1 = -2 y1 = -3 what are x2? y2?

OpenStudy (anonymous):

x is -2 and -3?

OpenStudy (anonymous):

x1 = -2 y1 = -3 what is x2? x2 = ??

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

Do you see the second point in your question? x2 =1 what is y2?

OpenStudy (anonymous):

-2 is xi and 1 is x2? then -3 is y1? and 1 is y2?

OpenStudy (anonymous):

Yes, now we need the slope formula x1 = -2 y1 = -3 x2 = 1 y2 = 1 slope = (y2 - y1)/(x2 - x1) so I will substitute the y's slope = (1 - -3)/(x2 - x1) or slope = 4/(x2 - x1) Can you substitute the x's?

OpenStudy (anonymous):

1.5?

OpenStudy (anonymous):

What did you get for x2 - x1? x2 = 1 x1 = -2 x2 - x1 =??

OpenStudy (anonymous):

x2 - x1 = 1 - -2 = 3 I got 4 and you got 3; therefore the slope is 4/3 <---------- answer